# Linear Bandits with Partially Observable Features

### v1

@@ C.1 Proof of Theorem 1
$=\mathbb{P}(E_{\widetilde{\theta}}|E_{\hat{\theta}})\cdot\mathbb{P}(E_{\hat{\theta}})+\mathbb{P}(E_{\widetilde{\theta}}|E_{\hat{\theta}}^{c})\cdot\mathbb{P}(E_{\hat{\theta}}^{c})$
$=\mathbb{P}(\widetilde{\theta}_{t}\geq 0|\hat{\theta}_{t}\geq 0)\cdot\mathbb{P}(\hat{\theta}_{t}\geq 0)+\mathbb{P}(\widetilde{\theta}_{t}\geq 0|\hat{\theta}_{t}<0)\cdot\mathbb{P}(\hat{\theta}_{t}<0)$
$\leq\exp\left(-\frac{t}{128\sigma^{2}}\right)+\mathbb{P}(\widetilde{\theta}_{t}\geq 0|\hat{\theta}_{t}<0).$ (20)
@@
Note that the total decision horizon $T>256\sigma^{2}\log(1/\delta^{\prime})$. For any $t>128\sigma^{2}\log(1/\delta^{\prime})$, we have $\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq 1/2-\delta^{\prime}$ This implies that for the rounds more than $T/2$, the suboptimal arm is expected to be played at least $(1/2-\delta^{\prime})T/2$ times for any $\delta^{\prime}\in(0,1/2)$, thus incurring
@@
For OFUL, we also present another analysis that requires no assumption such that the suboptimal arm is played for initial $t$ rounds, which is taken in Theorem 2 of Ghosh et al. [2017]. The optimal arm, arm 2, is selected when
@@
$\mathbb{P}\left(\widehat{\theta}_{t}>1/\sqrt{2t}\right)$ $\leq\mathbb{P}\left(-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{\sqrt{2t}}\right)$
@@ C.2 Proof of Theorem 2
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\sqrt{\sum_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|^{2}}=t^{-1/2}\|\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star}\|_{\mathbf{V}_{t}}$
Recall that $\widehat{\mathbf{w}}_{t}:=\mathbf{V}_{t}^{1/2}\widehat{\boldsymbol{\mu}}_{t}$ and $\mathbf{w}_{t}:=\mathbf{V}_{t}^{1/2}\boldsymbol{\mu}_{\star}$. Then
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}}^{\top}(\widehat{\boldsymbol{\mu}}_{t}-{\boldsymbol{\mu}}_{\star})|\leq t^{-1/2}\|\widehat{\mathbf{w}}_{t}-\mathbf{w}_{t}\|_{2}.$
To use Lemma 3, we prove a bound for
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-{\widetilde{\mathbf{x}}_{a}}^{\top}\mathbf{V}_{t}^{-1/2}{\mathbf{w}}_{t})\mathbf{V}_{t}^{-1/2}{\widetilde{\mathbf{x}}_{a}}\right\|_{\infty}.$
Let $\check{\mathbf{w}}_{t}^{L}:=\mathbf{V}_{t}^{1/2}\check{\boldsymbol{\mu}}_{t}^{L}$. By definition of $\widetilde{y}_{a,\tau}$,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t})\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}.$
$=\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)+\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t}\right)\mathbf{V}_{t}^{-1/2}\widetilde{\mathbf{x}}_{a}\right\|_{\infty}$
$\leq\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}$
$\quad+\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$
@@
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{w}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$ $=\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\mathbf{\widetilde{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$
$=\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}.$
@@
$\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}=\frac{1}{p}\max_{a\in[K]}|{\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}}|$
@@
$\max_{a\in[K]}|{\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}}|$ $\leq\max_{a\in[K]}\sigma\sqrt{2\sum_{\tau=1}^{t}\left(\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right)^{2}\log\frac{2Kt^{2}}{\delta}}$
$=\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\right)\mathbf{V}_{t}^{-1/2}\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
$\leq\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
$=\sigma\sqrt{2\log\frac{2Kt^{2}}{\delta}},$
and thus,
$\frac{1}{\sqrt{p}}\left\|\sum_{\tau=1}^{t}\epsilon_{a,\tau}\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}\leq\sigma\sqrt{\frac{2}{p}\log\frac{2Kt^{2}}{\delta}}$ (22)
Let $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$. Then the first term,
$\begin{split}&\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{V}_{t}^{-1/2}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}\\ &=\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}.\end{split}$ (23)
@@
$\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}=$ $\max_{a\in[K]}|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)|$
$\leq$ $\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\!\!\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{2}.$
Because $\widehat{\mathbf{w}}_{t}$ is a minimizer of Eq. 8, by Lemma 3 and Eq. 22,
@@
$\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{\mathbf{V}_{t}^{-1/2}\frac{1}{p}\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\mathbf{V}_{t}^{-1/2}}\leq 4\sigma\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p\lambda_{\min}\left(\mathbf{V}_{t}^{-1/2}\frac{1}{p}\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\mathbf{V}_{t}^{-1/2}\right)}}.$
Because $\phi_{a_{\tau},\tau}=p$ and the coupling event $\cap_{\tau\geq 1}\mathcal{M}_{\tau}$ holds with probability at least $1-\delta/t^{2}$,
$\sum_{\tau=1}^{t}\frac{1}{p}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}$ $=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{p}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$
$=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{\phi_{a,\tau}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$
$=\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$
$:=\mathbf{A}_{t}.$
Thus, under the coupling event $\cap_{\tau=1}^{t}\mathcal{M}_{\tau}$,
$\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}}\leq 4\sigma\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p\lambda_{\min}\left(\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right)}}.$
@@
By Corollary 1, with $\epsilon\in(0,1)$ to be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log\frac{2dt^{2}}{\delta}$, with probability at least $1-\delta/t^{2}$,
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon,$ (24)
which implies, $(1-\epsilon)\mathbf{I}_{K}\preceq\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}$ Thus,
@@
$\left\|\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right\|_{2}\leq\frac{4\sigma}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}.$
Now Eq. 23 is bounded by,
@@
$\begin{split}&\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\check{\mathbf{w}}_{t}^{L}-\mathbf{w}_{t}\right)\right\|_{\infty}\\ &\leq\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\frac{4\sigma}{1-\epsilon}\sqrt{\frac{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}{p}}.\end{split}$ (25)
With simple algebra,
$\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$=\max_{a\in[K]}\sqrt{\lambda_{\max}\left(\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\left(\mathbf{e}_{a}\mathbf{e}_{a}^{\top}\right)\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right)}$
$\leq\max_{a\in[K]}\sqrt{\lambda_{\max}\left(\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\mathbf{I}_{K}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right)}$
$=\left\|\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$=\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon$

### v2

@@ E.1 Proof of Theorem 1
$=\mathbb{P}(E_{\widetilde{\theta}}\mid E_{\hat{\theta}})\cdot\mathbb{P}(E_{\hat{\theta}})+\mathbb{P}(E_{\widetilde{\theta}}\mid E_{\hat{\theta}}^{c})\cdot\mathbb{P}(E_{\hat{\theta}}^{c})$
$=\mathbb{P}(\widetilde{\theta}_{t}\geq 0\mid\widehat{\theta}_{t}\geq 0)\cdot\mathbb{P}(\widehat{\theta}_{t}\geq 0)+\mathbb{P}(\widetilde{\theta}_{t}\geq 0\mid\widehat{\theta}_{t}<0)\cdot\mathbb{P}(\widehat{\theta}_{t}<0)$
$\leq\exp\left(-\frac{t}{128\sigma^{2}}\right)+\mathbb{P}(\widetilde{\theta}_{t}\geq 0\mid\widehat{\theta}_{t}<0).$ (18)
@@
Note that the total decision horizon $T>256\sigma^{2}\log(1/\delta^{\prime})$. For any $t>128\sigma^{2}\log(1/\delta^{\prime})$, we have $\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq 1/2-\delta^{\prime}$. This implies that for the rounds more than $T/2$, the suboptimal arm is expected to be played at least $(1/2-\delta^{\prime})T/2$ times for any $\delta^{\prime}\in(0,1/2)$, thus incurring
@@
For OFUL, we also present another analysis that requires no assumption such that the suboptimal arm is played for initial $t$ rounds, which is taken in Theorem 2 of Ghosh et al. (2017). The optimal arm, arm 2, is selected when
@@
$\mathbb{P}\left(\widehat{\theta}_{t}>\frac{1}{\sqrt{2t}}\right)$ $\leq\mathbb{P}\left(-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{\sqrt{2t}}\right)$
@@ E.2 Proof of Theorem 2
$\max_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{L}_{t}-{\boldsymbol{\mu}}_{\star})|\leq\sqrt{\sum_{a\in[K]}|\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}^{L}_{t}-{\boldsymbol{\mu}}_{\star})|^{2}}=t^{-1/2}\|\widehat{\boldsymbol{\mu}}^{L}_{t}-{\boldsymbol{\mu}}_{\star}\|_{\mathbf{V}_{t}}.$
To use Lemma 3, we prove a bound for $\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{t}\right)\widetilde{\mathbf{x}}_{a}\|_{\infty}$. By definition of $\widetilde{y}_{a,\tau}$,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(\widetilde{y}_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\widetilde{\mathbf{x}}_{a}\right\|_{\infty}$
$=\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)+\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\widetilde{\mathbf{x}}_{a}\right\|_{\infty}$
$\leq\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}+\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{\widetilde{x}}_{a}\right\|_{\infty}.$
@@
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\widetilde{\mathbf{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$ $=\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(a_{\tau}=a)}{\phi_{a,\tau}}\left(y_{a,\tau}-\mathbf{\widetilde{x}}_{a}^{\top}\boldsymbol{\mu}_{\star}\right)\mathbf{\widetilde{x}}_{a}\right\|_{\infty}$
$=\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{\tau}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}.$
@@
$\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{\tau}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}=\frac{1}{p}\max_{a\in[K]}\left|\sum_{\tau=1}^{t}\epsilon_{\tau}\mathbf{e}_{a}^{\top}\mathbf{\widetilde{x}}_{a_{\tau}}\right|.$
@@
$\max_{a\in[K]}\left|\sum_{\tau=1}^{t}\epsilon_{\tau}\mathbf{e}_{a}^{\top}\mathbf{\widetilde{x}}_{a_{\tau}}\right|$ $\leq\max_{a\in[K]}\sigma\sqrt{2\sum_{\tau=1}^{t}\left(\mathbf{e}_{a}^{\top}\mathbf{\widetilde{x}}_{a_{\tau}}\right)^{2}\log\frac{2Kt^{2}}{\delta}}$
$=\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\left(\sum_{\tau=1}^{t}\mathbf{\widetilde{x}}_{a_{\tau}}\mathbf{\widetilde{x}}_{a_{\tau}}^{\top}\right)\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
$\leq\max_{a\in[K]}\sigma\sqrt{2\mathbf{e}_{a}^{\top}\left(\mathbf{V}_{t}\right)\mathbf{e}_{a}\log\frac{2Kt^{2}}{\delta}}$
$=\sigma\widetilde{\sigma}_{\max}\sqrt{2t\log\frac{2Kt^{2}}{\delta}},$
where the last equality follows by the definition $\widetilde{\sigma}^{2}_{\max}=\max_{a\in[K]}\mathbf{e}_{a}^{\top}(\sum_{a\in[K]}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top})\mathbf{e}_{a}$. Thus,
$\frac{1}{p}\left\|\sum_{\tau=1}^{t}\epsilon_{\tau}\mathbf{\widetilde{x}}_{a_{\tau}}\right\|_{\infty}\leq\frac{\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{2t\log\frac{2Kt^{2}}{\delta}}.$ (20)
Let $\mathbf{A}_{t}:=\sum_{\tau=1}^{t}\sum_{a\in[K]}\mathbb{I}(\widetilde{a}_{\tau}=a)\phi_{a,\tau}^{-1}\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}$. Then the first term,
$\left\|\sum_{\tau=1}^{t}\sum_{a\in[K]}\left(1-\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\right)\mathbf{\widetilde{x}}_{a}\mathbf{\widetilde{x}}_{a}^{\top}\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}=\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}.$ (21)
@@
$\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}=$ $\max_{a\in[K]}|\mathbf{e}_{a}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)|$
$\leq$ $\max_{a\in[K]}\left\|\mathbf{e}_{a}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1/2}\right\|_{2}\!\!\left\|\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right\|_{\mathbf{A}_{t}}.$
Because $\check{\boldsymbol{\mu}}_{t}^{L}$ is a minimizer of Eq. 8, by Lemma 3 and Eq. 20,
@@
$\left\|\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right\|_{\mathbf{A}_{t}}\leq 4\sigma\widetilde{\sigma}_{\max}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{p\lambda_{\min}\left(\mathbf{A}_{t}\right)}}.$
@@
By Corollary 1, with $\epsilon\in(0,1)$ to be determined later, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-\delta/t^{2}$,
$\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}\leq\epsilon.$ (22)
Eq. 22 implies $(1-\epsilon)\mathbf{I}_{K}\preceq\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}$ and $(1-\epsilon)\mathbf{V}_{t}\preceq\mathbf{A}_{t}$. Thus,
@@
$\left\|\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right\|_{\mathbf{A}_{t}}\leq\frac{4\sigma\widetilde{\sigma}_{\max}}{p}\sqrt{\frac{2t(d+d_{h})\log(2Kt^{2}/\delta)}{(1-\epsilon)\lambda_{\min}\left(\mathbf{V}_{t}\right)}}=\frac{4\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})\log(2Kt^{2}/\delta)}{1-\epsilon}},$
and Eq. 21 is bounded by,
@@
$\begin{split}\left\|\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{L}-\boldsymbol{\mu}_{\star}\right)\right\|_{\infty}\ &\leq\max_{i\in[K]}\left\|e_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1/2}\right\|_{2}\frac{4\sigma\widetilde{\sigma}_{\max}}{p\widetilde{\sigma}_{\min}}\sqrt{\frac{2(d+d_{h})\log(2Kt^{2}/\delta)}{1-\epsilon}}\\ &\leq\max_{i\in[K]}\left\|e_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}\frac{4\sigma\widetilde{\sigma}_{\max}}{p(1-\epsilon)\widetilde{\sigma}_{\min}}\sqrt{2(d+d_{h})\log\frac{2Kt^{2}}{\delta}}.\end{split}$ (23)
By Eq. 22,
$\max_{i\in[K]}\left\|\mathbf{e}_{i}^{\top}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}$ $=\max_{i\in[K]}\left\|\mathbf{e}_{i}^{\top}\mathbf{V}_{t}^{1/2}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$\leq\widetilde{\sigma}_{\max}\sqrt{t}\left\|\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\right\|_{2}$
$\leq\widetilde{\sigma}_{\max}\sqrt{t}\epsilon.$